OpenStudy (schoolgeek15):

Two equations are shown: • Equation 1: 3 4 (x−12)=12 •Equation 2: 3 4 y−12=12 Solve each equation. Then, enter a number in each box to make this statement true. The value of x is: The value of y is:

OpenStudy (schoolgeek15):

@mangotangochick Help me?

OpenStudy (mangotangochick):

I don't know what the 3 and 4 numbers signify xD but the question can be done nonetheless: 1) (x-12)=12 The brackets are there just to throw you off, they are not needed, so we can remove the brackets/expand the left hand side of this equation: x-12 = 12 therefore if we add 12 to both sides, we end up with: x = 0 2) y-12 = 12 so we also have to add 12 to both sides (to get y on its own) so that: y = 0 We end up with x = 0 and y = 0 (x = y)

OpenStudy (mangotangochick):

Wait, 12 + 12 = 24 so x and y are both 24 IM SUCH A DUMMY SORRY

OpenStudy (mangotangochick):

I wish there was an edit button on this site

OpenStudy (mangotangochick):

Please see this before you get something wrong on your homework!!!!!!

OpenStudy (radar):

You may have this closed too early. I too don't understand the significance of the 3 and 4, but the equations can be solved by inspection. 1. x - 12 = 12 x must be 24 same as y - 12 = 12. I see that mangotangochick has returned with a correction